A stock had returns of 16.59 percent (1 year ago), -27.64 percent (2 years ago), X (3 years ago), and 23.57 percent (4 years ago) in each of the past 4 years. Over the past 4 years, the geometric average annual return for the stock was 4.4 percent. What was the arithmetic average annual return for the stock over the past 4 years? Answer as a rate in decimal format so that 12.34% would be entered as .1234 and 0.98% would be entered as .0098.
Answer #1

r1 = 16.59%

r2 = -27.64%

r3 = X

r4 = 23.57%

Formula for Geometric Maen:-

Geometric Mean= \sqrt[n]{(1+r_1)*(1+r_2)*...(1+r_n)}-1

0.044= \sqrt[n]{(1+0.1559)*(1-0.2764)*(1+X)*(1+0.2357)}-1

1.044= \sqrt[n]{(1.1559)*(0.7236)*(1+X)*(1.2357)}

Taking 4-Square on both sides,

1.18796048409= 1.04249242306*(1+X)}

1.18796048409= 1.04249242306+1.04249242306*X

X = 13.9539%

So, Return in year 3 is 13.9539%

Now, Calculating Arithmetic Average Annual return:-

Arithenetic Average Annual return = (r1 + r2 + r3 + r4)*/4

= (16.59% - 27.64% + 13.9539% + 23.57%)/4

= 6.62%

So, Arithmetic Average Annual return is 0.0662